15-th Conference of the International Linear Algebra Society

April 28, 2008

Professor Dr. Zeyad Abdel Aziz Al Zhour

Zarqa Private University, Zarqa, Jordan

Dear Professor Al Zhour:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Matrix results on weighted Drazin inverse and some applications” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx